The length of a rectangle is three more than triple the width. If the perimeter is 94 inches, find the dimensions
Answer by greenestamps(13200)



A good easy problem for solving using basic algebra....

Let x be the width.
Then, per the given information, the length is 3x+3. ("three more than triple the width")

The perimeter (twice width plus twice length) is 94 inches, so the length plus the width is 94/2 = 47 inches.

Using the expressions for the width and length, write and solve an equation that says length plus width is 47 inches:

(x)+(3x+3) = 47
4x+3 = 47
4x = 47-3 = 44
x = 44/4 = 11

ANSWERS:
width = x = 11
length = 3x+3 = 3(11)+3 = 33+3 = 36